What is the Fahrenheit temperature of a person with a 39.3ºC fever?
Answer #1

The temperature T in degrees Fahrenheit (°F) is equal to the temperature T in degrees Celsius (°C) times 9/5 plus 32:

T(°F) = T(°C) × 9/5 + 32 = 39.3 × 9/5 + 32= 102.74 °F